The table shows a bank balance sheet. The desired ratio on the deposits is 5 %. There is no currency drain. What is the quantity of loans and the quantity of total deposit when the bank has no excess reverses.

reverses at the Fed= 25 checkabledeposits=105
cash in vault=10
securities= 60 saving deposits=95
loans= 105
1
Expert's answer
2016-10-20T12:11:11-0400
If the bank has no excess reverses, then the quantity of loans is L = reserves + cash in vault + securities + loans = 25 + 10 + 60 + 105 = 200 and the quantity of total deposit is D = checkable deposits + saving deposits = 105 + 95 = 200.